OpenStudy (anonymous):

9x -5y +16 =0 Having trouble finding the slope?

OpenStudy (anonymous):

First, you want to get that into slope-intercept form by getting that y by itself.

OpenStudy (callisto):

or for a equation having general form ax+by+c=0 slope = -a/b x-int = -c/a y-int = -c/b

OpenStudy (anonymous):

Change that into y=mx + c form and here m is the slope

OpenStudy (anonymous):

bring y on the left side....only y no coffecient or anything else

OpenStudy (anonymous):

and the coefficient of x is the slope

OpenStudy (anonymous):